# Neural Audio Codecs for Prompt-Driven Universal Source Separation

### v1

@@ 3.3 Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
Dimensionality reduction and compression.
For a typical 1 s audio, 32 kHz STFT used by spectrogram-domain baselines, the mixture spectrogram $X\in\mathbb{C}^{F\times T}$ (e.g., $F=2\cdot 1024$, $T\approx 100$) contains about $|X|\approx 204{,}800$ complex coefficients. In contrast, a neural audio codec latent $Z\in\mathbb{R}^{d\times T_{c}}$ is far more compact: at 16 kHz with $d=64$ and $T_{c}=\tfrac{16000}{320}\approx 50$, |Z|=3,200|Z|=3{,}200; at 32 kHz with $T_{c}=\tfrac{32000}{320}\approx 100$, |Z|=6,400|Z|=6{,}400. Thus, the ratio $|Z|/|X|$ is $\approx 1.5625\times 10^{-2}$ for 16 kHz and $\approx 3.125\times 10^{-2}$ for 32 kHz , implying a $\sim 32$$64\times$ reduction in representation size. Operating directly in $Z$ therefore cuts memory bandwidth and MACs for the separator $Mask(\cdot)$ while retaining perceptually salient structure imparted by the codec, making latent-domain separation both compute-efficient and semantically aligned. Moreover, the compact channel dimension ($d=64$) enables a smaller masker (fewer parameters in $Q/K/V$ and MLP projections) and more effective self-attention.
3.3.1 Why NAC latents (vs. spectrograms) and how the codec prior enables separation?
$Enc(.):x\mapsto Z\in\mathbb{R}^{d\times T_{c}}.$
$\tilde{Z}_{k}\;=\;M_{k}\odot Z,\qquad\tilde{y}_{k}\;=\;Dec(\tilde{Z}_{k}),$
$Mask(.,.):\;Z\in\mathcal{M}_{\text{latent}}\;\to\;\hat{Z}_{k},$
$g(.,.):\;X\in\mathbb{C}^{F\times T}\;\to\;\hat{Y}_{k},$
Hierarchical (RVQ) structure that benefits separation. The codec applies Residual Vector Quantization (RVQ) to $Z$, producing discrete codes $A=[a_{t}\in[K]^{N_{q}}]_{t=1}^{T_{c}}$ with $K=1024$ and $N_{q}$ quantizers. Codebook lookup yields embeddings
The overall loss is
3.3.2 Transformer-based masker vs. decoder-style generation (specific design choice).
3.3.3 Why We Integrated FiLM Conditioning into NAC-Based Separation
We deliberately integrate Feature-wise Linear Modulation (FiLM) Perez et al. (2018) inside the transformer masker to inject text semantics while preserving the codec manifold and keeping $Enc(\cdot)$$Dec(\cdot)$ frozen.
Targeted placement (masker, mid-layers).
Lightweight computation (overhead and parameterization).
Non-iterative inference (single forward pass).
Manifold preservation and stability.
3.3.4 Why report results on $Z$ (continuous latents) and how to extend to code streams $A\!\to\!E$
Why we evaluate on continuous latents $Z$.
We choose to perform—and therefore report—separation on the continuous DAC encoder latents $Z=Enc(x)\in\mathbb{R}^{d\times T_{c}}$ for the following concrete reasons:
  1. 1.
  2. 2.
  3. 3.
How to extend the same model to discrete code streams.
$A$ $=[a_{t}\in[1024]^{N_{q}}\mid t\in[T_{c}]],\qquad e_{t}\;=\;\sum_{i=1}^{N_{q}}\mathrm{lookup}\!\big{(}a_{t}^{(i)}\big{)},$ (5)
$E$ $=[e_{t}]_{t=1}^{T_{c}}\;\approx\;Z,\qquad\tilde{E}_{s}\;=\;M_{s}\odot E,\qquad\tilde{y}_{s}(t)\;=\;Dec(\tilde{E}_{s})$ (6)
$\tilde{E}_{s}\;=\;M_{s}\odot E\;\approx\;M_{s}\odot Z\;=\;Z_{s}$
Embedding alignment (CodecFormer-EL Yip et al. (2024a)). To tighten $\tilde{E}_{s}\approx Z_{s}$, we can optimize an embedding-level consistency loss as in CodecFormer-EL Yip et al. (2024a)instead of our SI-SDR objective:
$\mathcal{L}_{\mathrm{emb}}\;=\;\sum_{s}\|\tilde{E}_{s}-Z_{s}\|_{1}$
3.3.5 Deployment advantages of CodecSep over spectrogram-domain separators (AudioSep)
Pipeline comparison (server-side separation).
Traditional (spectrogram) pipeline:
CodecSep pipeline:
Complexity accounting.
In the common (code-stream) case, CodecSep removes both decode and re-encode on the server. Moreover, within the separator, CodecSep operates on $Z/E$ with $|Z|\ll|X|$ (cf. Sec. 3.3), reducing activation memory and bandwidth throughout attention and MLP blocks.
Interface compatibility with codec bitstreams.
Why spectrogram-domain systems incur extra overhead.
Operational advantages of CodecSep.
  • Eliminates redundant codec cycles in server workflows: With code streams, we reconstruct embeddings $E\approx Z$ by codebook lookup (Sec. 3.3.4), avoiding server-side decode/re-encode; only the edge decodes the final stems.
@@ 4 Experiments
4.1 Datasets
4.1.1 Divide and Remaster v 2.0 (dnr-v2)
dnr-v2 Petermann et al. (2022) dataset consists of $60s$-duration artificial mixtures of speech, music, and SFX sampled from LibriSpeech Panayotov et al. (2015), Free Music Archive (FMA) Defferrard et al. (2017), and Freesound Dataset $50K$ (FSD50K50K) Fonseca et al. (2022), respectively. It includes $3,406$ ($56.7hrs$) training, $487$ ($8.13hrs$) validation, and $973$ ($16.22hrs$) test mixtures, each provided with its three individual source audios. The mixtures are generated by normalizing each source to fixed Loudness Units Full-Scale (LUFS) levels: 17-17 dB (speech), 24-24 dB (music), and 21-21 dB (SFX), with $\pm 2$ dB random perturbations. Any source exceeding a peak threshold is normalized to $0.5$ dB. The sources are mixed and normalized to 27-27 dB LUFS with additional random perturbations. The validation and test sets are trimmed for silence and split into $5s$ or $10s$ segments. Segments where sources are present for less than $50\%$ of the duration are removed, resulting in $2,852$ ($\approx 3.96hrs$) validation and $1,840$ ($\approx 5.11hrs$) test mixtures.
4.1.2 Open-Domain Benchmarks
We benchmark on five open-domain datasets spanning captioned audio, environmental sounds, and large multi-event corpora: AudioCaps Kim et al. (2019), an AudioSet-derived collection of >46>46k $10$ s YouTube clips paired with human-written captions describing the dominant sound events (used by us to synthesize training and test mixtures); ESC-50 Piczak , a curated environmental sound dataset of 2,0002{,}000 clips ($5$ s each) organized into $50$ classes with $40$ examples per class across five meta-categories (animals, natural, human non-speech, domestic, exterior/urban); Clotho-v2 Drossos et al. (2020), 6,9746{,}974 audio samples (15153030 s) each annotated with five human captions (8–20 words) covering open-domain events; AudioSet-eval Gemmeke et al. (2017), the evaluation split of AudioSet comprising human-labeled $10$ s YouTube clips over an ontology of $632$ audio event classes in a multi-label setting; and VGGSound Chen et al. (2020a), an AudioSet-derived audio–visual corpus with 550+550{+} hours of $10$ s segments covering a wide variety of everyday sound categories. For AudioCaps we form both training and testing mixtures by summing three clips (validation segmented into $5$ s, test preserves clips up to $20$ s), while for ESC-50, Clotho-v2, AudioSet-eval, and VGGSound we construct test-only mixtures using the same three-clip protocol.
4.2 Evaluation
We evaluate CodecSep alongside TDANet Li et al. (2023); Pons et al. (2024), CodecFormer Yip et al. (2024b), SDCodec Bie et al. (2024), and AudioSep Liu et al. (2024) on universal source separation conducted on the dnr-v2 and AudioCaps test splits. We report the scale-invariant signal-to-distortion ratio (SI-SDR) Luo and Mesgarani (2019); Le Roux et al. (2019) for signal quality assessment and Virtual Speech Quality Objective Listener (ViSQOL) Chinen et al. (2020) for perceptual quality assessment. ViSQOL Chinen et al. (2020) evaluates the perceptual similarity between the ground truth $x$ and the estimated signal $\tilde{x}$ using a spectro-temporal approach. It then predicts a Mean Opinion Score - Listening Quality Objective (MOS-LQO), which ranges from $1$ (lowest quality) to $5$ (highest quality). In addition, we evaluate the inference efficiency of our models in a hardware-agnostic setting by reporting Multiply-and-Accumulate operations (MACs) using Torchprofile  [^1] https://github.com/zhijian-liu/torchprofile.git averaged over 8 audios of $2s$ duration each.
4.3 Training
The complete model, including the query module $query(.)$, is trained for $400K$ iterations with DAC Kumar et al. (2023) and CLAP Wu et al. (2023) modules frozen. Validation is conducted every $5K$ iterations and test every $10K$ iterations. We use ADAM Kingma and Ba (2017) as our optimizer and train with a batch size of $4$ examples, each $2$ seconds in duration, and a learning rate of $1.5e^{-4}$ on a single 2424GB NVIDIA A-$30$ GPU. Training employs a $ReduceLRonPlateau$ Mukherjee et al. (2019) scheduler, which reduces the learning rate by a factor of $0.5$ if the validation loss does not improve for two consecutive validation checks. We train two versions of CodecSep, one using the dnr-v2 dataset and the other using AudioCaps, to evaluate performance across different training distributions. We refer to these models using the prefixes +dnr-v2 and +AudioCaps, respectively, to indicate which dataset each model was trained on.
@@ 4.3 Training
Model Metric ($\uparrow$) Separation
Music Speech Sfx
@@
(zero-shot) ViSQOL $2.36^{\pm 0.71}$ $2.43^{\pm 0.70}$ $2.15^{\pm 0.79}$
(zero-shot) ViSQOL $2.27^{\pm 0.53}$ $2.56^{\pm 0.47}$ $2.06^{\pm 0.72}$
@@ 4.4 Results and Discussions
Tables 16 present universal source separation results, prompt–granularity analyses, architectural ablations, cross-dataset generalization, paraphrase robustness, and full inference complexity under matched training/evaluation protocols. Table 1 reports dnr-v2 test results for speech, music, and SFX: text-guided models use generic prompts for speech/music and ground-truth compositional captions for SFX; we also include a masker ablation to isolate the role of the Transformer masker. Table 2 studies SFX prompt granularity across three regimes—(i) fixed-stem, non-text baselines (TDANet, CodecFormer, SDCodec), (ii) generic 3-stem prompts ({music, speech, sfx}), and (iii) a universal setup with fine-grained, compositional SFX prompts—thereby aligning input conditions for fair comparison with fixed-head systems. Table 3 isolates architecture: decoder-style generation (CodecFormer) vs. an unguided 3-stem masker variant and its text-guided counterpart, all operating in the codec latent space. To assess out-of-domain generalization, Table 4 evaluates on AudioCaps-test with mixtures of three randomly sampled sources and prompts drawn from captions (not tied to fixed labels). Beyond AudioCaps/dnr-v2, Table 5 benchmarks on four additional open-domain corpora (ESC-50, Clotho-v2, AudioSet-eval, VGGSound) under the same prompt protocol. Table 6 probes prompt paraphrasing robustness by replacing the generic speech/music prompts with unseen synonymic variants at inference (zero-shot paraphrase test). Finally, Table 7 compares end-to-end and architecture-only GMACs for spectrogram-domain separation versus codec-latent masking, including the practical code-stream case. All models are evaluated against the original (uncompressed) ground truth; our methods are highlighted in bold; and we report mean and standard deviation ($1\sigma$) over each evaluation set.
Source Separation Performance on dnr-v2 (Table 1). CodecSep+dnr-v2 outperforms both pretrained AudioSep (zero-shot scenario) and retrained AudioSep+dnr-v2 across all categories, with significant SI-SDR gains in speech ($\mathbf{9.97}$ vs. 4.924.92/ $7.68$ dB), music ($\mathbf{1.15}$ vs. -2.462.46/ -$5.55$ dB), and SFX ($\mathbf{0.89}$ vs. -0.340.34/ -$4.66$ dB). Regarding ViSQOL, CodecSep matches or exceeds AudioSep in speech and music but slightly underperforms in SFX. While AudioSep is trained on a diverse dataset with sfx prompts covering isolated events and complex mixtures, CodecSep is trained exclusively on long-form, compositional SFX prompts from dnr-v2. Despite this narrower prompt exposure, CodecSep remains robust and achieves the highest overall SI-SDR. As part of our ablation study, we evaluate a lightweight variant—CodecSep+dnr-v2 (ablate Masker)—that removes the transformer masker and applies text conditioning directly to the encoder. While it performs comparable to AudioSep+dnr-v2 in SI-SDR, it achieves better perceptual quality for speech. The drop in separation quality stems from FiLM distorting the encoder’s latent space, which holds the mixture representation.
@@
Effect of SFX Prompt Granularity during training (Table 2). We evaluate three regimes: (i) fixed-stem baselines without text guidance (TDANet, CodecFormer, SDCodec), (ii) generic 3-stem prompts (“music/speech/sfx”), and (iii) universal prompting that retains generic cues for speech/music but uses fine-grained, compositional SFX descriptions. Across settings, CodecSep is competitive with or stronger than fixed-stem baselines on SFX while maintaining comparable speech/music quality, despite those baselines being explicitly supervised per head. Under matched generic-prompt training/evaluation, CodecSep remains robust and surpasses the spectrogram-domain AudioSep, indicating that effective separation does not hinge on carefully crafted prompts. Moreover, replacing the single “sfx” label with detailed SFX prompts consistently sharpens SFX extraction and, importantly, improves speech and music stems as well—suggesting that richer SFX supervision enhances overall scene disentanglement. While these controlled studies cover multiple prompt granularities, we expect additional gains from training on larger, more diverse corpora with a spectrum of prompt specificities, which we leave for future work.
@@
Model Metric ($\uparrow$) Separation
Music Speech Sfx
@@
Universal: “music”,“speech”, fine-grained sfx prompt
AudioSep SI-SDR $2.46^{\pm 4.06}$ $4.92^{\pm 4.21}$ $0.34^{\pm 5.39}$
AudioSep + dnr-v2 SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$ $4.66^{\pm 3.68}$
ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$ $2.32^{\pm 0.7}$
ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
Why use a Transformer masker instead of a decoder?(Table 3). We compare (i) CodecFormer (decoder-style source generation), (ii) CodecSep (unguided, 3-stem) which uses the CodecFormer Transformer as a masker over codec latents, and (iii) CodecSep (text-guided). The results on dnr-v2-test exhibit two clear trends. First, replacing decoder-style generation with masking consistently strengthens separation across music, speech, and SFX. This aligns with our design rationale: in the DAC latent domain, the masker modulates existing, semantically structured content instead of synthesizing new signals, which (a) reduces artifacts and cross-talk leakage, (b) preserves long-range periodicity/timbre and transient organization already encoded by the codec, and (c) stabilizes optimization compared to end-to-end generation. Second, adding text guidance yields a further uniform improvement. The masker formulation concentrates Transformer capacity on selection (“where/how much” to pass) rather than generation (“what” to produce). Taken together, these findings support our design choice: in the codec latent space, a FiLM-conditioned Transformer masker is a more effective and robust allocator of model capacity than a decoder, with text guidance further sharpening that allocation.
@@
Model Metric ($\uparrow$) Separation
Music Speech Sfx

### v2

@@ 3.2 Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
with frozen DAC backbone $Enc(\cdot),Dec(\cdot)$ and a FiLM-conditioned transformer masker $Mask(\cdot,e_{\tau})$ that estimates $M_{s}\!\in\![0,1]^{d\times T}$.
Why NAC latents vs. spectrograms.
Operating on NAC latents $Z$ slashes dimensionality while preserving perceptual factors. For $1\,\mathrm{s}$ audio at $32\,\mathrm{kHz}$, complex STFT with N=1024N{=}1024 and hop size M=320M{=}320 samples has $T_{\mathrm{spec}}\!\approx\!100$ frames and $F=2\times 1024$ (Re+Im) scalars per frame, so $F\cdot T_{spec}\!\approx\!204{,}800$. A $16\,\mathrm{kHz}$ DAC with width d=64d{=}64 and the same $M$ yields $T\!\approx\!50$ and $d\cdot T{=}64\times 50{=}3{,}200$ ($\sim 64{\times}$ smaller), shrinking $Q/K/V$ and MLP sizes and easing self-attention. Similarly, for $32\,\mathrm{kHz}$ NACs like EnCodec, $T\!\approx\!50$ with d=128d{=}128, so attention/MLPs still operate on $\sim 32{\times}$ smaller latents than complex STFTs. Crucially, $Enc(\cdot)$ organizes $Z$ on a discriminative, perceptually aligned manifold , making selection (masking) easier than representation learning from raw $X$. In spectrogram systems, the separator must first learn a high-level latent from $X$ (via CNN/UNet) and then separate, coupling abstraction and masking and inflating parameters/compute.
Masking over codec latents (leveraging the codec prior).
Because the DAC codec induces a strong semantic prior in its latent space via residual vector quantization (RVQ) and perceptual/adversarial training, we mask the codec latents rather than generate sources from scratch as in CodecFormer. RVQ creates a coarse-to-fine hierarchy in $Z{=}Enc(x)\in\mathbb{R}^{d\times T}$: early stages capture coarse structure (e.g., low-frequency content, timbre), while later stages refine residual detail (e.g., high-frequency components, transients) Wang et al. (2023). We exploit this organization with a FiLM-conditioned transformer masker that predicts a soft mask $M_{s}\in[0,1]^{d\times T}$ and applies it element-wise, yielding source latent estimate $\tilde{Z}_{s}=M_{s}\odot Z$. In contrast to learning a generator $Sep(Z,e_{\tau}):Z\!\to\!\tilde{Z}_{s}$ as in CodecFormer, learning a mask $Mask(Z,e_{\tau}):Z\!\to\!M_{s}$ on the compact, semantically organized codec manifold both exploits the codec prior more effectively and yields a more stable optimization that converges faster. Moreover, masking in the denoised, low-dimensional codec space is fundamentally easier than masking in the high-dimensional, noisy spectrogram domain. This selection-centric design (i) constrains learning to modulation of existing latent content, (ii) avoids hallucination and reduces leakage because no new signals are synthesized, and (iii) preserves long-horizon structure (periodicity, timbre, transients) already organized by the codec, yielding stable, low-artifact separations.
Why FiLM inside the masker.
Placing FiLM in the masker (not in $Enc/Dec$) targets the selection step, preserves the codec manifold, and adds negligible overhead (two vectors per layer). Conditioning is non-iterative (single forward pass), maintaining low latency for edge/server workflows.
Continuous latents $Z$ vs. discrete codes $A$ (and deployment path).
For training and analysis, we operate on continuous latents $Z{=}Enc(x)\in\mathbb{R}^{d\times T}$: (i) gradients flow cleanly through $Mask(\cdot,e_{\tau})$ and $Dec(\cdot)$ with a frozen codec (no straight‐through estimators), yielding stable convergence; (ii) RVQ pretraining regularizes $Z$ so pitch, timbre, onsets/transients, and textures are hierarchically organized, providing a richer, more disentangled signal for FiLM; and (iii) $Z$ avoids run‐to‐run variance from codebook utilization (e.g., late RVQ sensitivity, bitrate truncation), reducing the need for special regularizers. For deployments with compressed bitstreams, we reconstruct embeddings by codebook lookup and use the same masker:
$A$ $=[a_{t}\in[1024]^{N_{q}}\mid t\in[T]],\qquad e_{t}\;=\;\sum_{i=1}^{N_{q}}\mathrm{lookup}\!\big(a_{t}^{(i)}\big),$ (1)
$E$ $=[e_{t}]_{t=1}^{T}\;\approx\;Z,\qquad\tilde{E}_{s}\;=\;M_{s}\odot E,\qquad\tilde{y}_{s}(t)\;=\;Dec(\tilde{E}_{s})$ (2)
When a codes-out interface is desired, we re-quantize masked embeddings and optionally decode:
$\hat{A}_{s}=Quant(\tilde{E}_{s}),\qquad\hat{E}_{s}=\mathrm{lookup}(\hat{A}_{s}),\qquad\tilde{y}_{s}(t)=Dec(\hat{E}_{s}).$
By design, $E\!\approx\!Z$ at the operating bitrate and $Dec(E)$ already yields high-fidelity reconstructions; because our separator is a masker (selective modulation) rather than a generator, swapping $Z\!\to\!E$ preserves the semantics needed for separation with no architectural change. While we report results on $Z$ to isolate separator performance and maintain stable optimization, we also evaluate the bitstream path by feeding reconstructed embeddings $E$ (codes-in) to the same trained masker without any fine-tuning; performance remains competitive relative to the $Z$ path. The residual gap can be narrowed with light fine-tuning the masker on $E$ or optimizing an embedding-consistency loss (cf. CodecFormer-EL) in place of, or alongside, SI-SDR: $\mathcal{L}_{\mathrm{emb}}=\sum_{s}\|\tilde{E}_{s}-Z_{s}\|_{1}$. In deployment, the variant simply replaces the masker input with $E$ and optionally re-quantizes for codes-out.
Deployment advantages over spectrogram-domain separators.
In realistic pipelines, edge devices already run a codec and send code streams. Spectrogram models must decode $\to$ separate on $X$ $\to$ re-encode, adding latency and energy; CodecSep masks in the codec domain and can return separated code streams directly:
$\text{Edge: }Enc(\cdot)\Rightarrow\text{Code}\Rightarrow\text{Server: reconstruct }E\!\approx\!Z\ \&\ Mask(\cdot,e_{\tau})\Rightarrow\text{Code}\Rightarrow\text{Edge: }Dec(\cdot).$
Concretely, with codec costs $C_{\mathrm{enc}},C_{\mathrm{dec}}$, spectrogram separator cost $C_{\mathrm{spec}}$, and CodecSep masker cost $C_{\mathrm{mask}}$:
$\text{Code-stream input: }\ \text{AudioSep}=C_{\mathrm{dec}}{+}C_{\mathrm{spec}}{+}C_{\mathrm{enc}},\quad\text{CodecSep}=C_{\mathrm{mask}}.$
Within the separator, working on $Z/E$ (small $d,T$ with $|Z|\!\ll\!|X|$) shrinks attention/MLP activations, enabling tighter batching and lower memory bandwidth. Interface compatibility is immediate when only codes $A$ are available: perform lookup to $E$, apply the same FiLM-conditioned masker, and optionally re-quantize to emit $\hat{A}_{s}$. Overall, CodecSep (i) eliminates redundant decode/re-encode in server workflows; (ii) preserves the codec manifold (periodicity, timbre, transients) by modulating $Z/E$ instead of rebuilding $X$; and (iii) keeps conditioning non-iterative (single forward pass), yielding low-latency, high-fidelity stems at scale.
See Appendix A for a full rationale covering latent-domain masking vs. spectrogram-domain generation, the RVQ-induced semantic prior, FiLM placement inside the masker, continuous $Z$ vs. discrete $A{\to}E$ operation (and codes-in/out), and complexity/deployment comparisons.
@@ 4 Experiments
Datasets.
We evaluate across a controlled multi-stem corpus and multiple open-domain benchmarks. For in-domain experiments, we adapt Divide and Remaster v2 (dnr-v2) Petermann et al. (2022) from fixed-label three-stem separation to universal, prompt-driven separation by replacing source labels with natural-language queries. Speech and music are queried with broad category prompts (e.g., “speech”, “music”), while SFX stems are queried using long-form, compositional prompts (>1>1 overlapping sources) synthesized from FSD50K’s hierarchical annotations Fonseca et al. (2022), combining fine-grained classes with parent categories (e.g., “dog barking, Animal, engine rumbling, motor vehicle”). To assess cross-domain generalization, we form three-source mixtures on AudioCaps Kim et al. (2019) (used for both training and testing in our open-domain setting) and construct test-only three-source mixtures from ESC-50 Piczak , Clotho-v2 Drossos et al. (2020), AudioSet-eval Gemmeke et al. (2017), and VGGSound Chen et al. (2020a). Dataset construction details, clip durations, split statistics, and segmentation rules are provided in the Appendix B.
Evaluation.
We compare CodecSep against representative spectrogram- and codec-domain baselines—TDANet Li et al. (2023); Pons et al. (2024), CodecFormer Yip et al. (2024b), SDCodec Bie et al. (2024), and the text-guided spectrogram model AudioSep Liu et al. (2024). We report objective signal fidelity via scale-invariant signal-to-distortion ratio (SI-SDR) Luo and Mesgarani (2019); Le Roux et al. (2019) and perceptual quality via ViSQOL Chinen et al. (2020), which measures spectro–temporal similarity between the estimate $\tilde{x}$ and reference $x$ and maps it to a 1–5 MOS-LQO scale. Following prior work (e.g., SDCodec Bie et al. (2024)), we use ViSQOL as a proxy for perceptual listening quality and complement it with a human MOS-LQS study comparing real-world outputs from CodecSep (trained on dnr-v2) and the publicly released AudioSep. To quantify efficiency in a hardware-agnostic manner, we report multiply–accumulate operations (MACs) using torchprofile [^1] https://github.com/zhijian-liu/torchprofile under matched input durations (2 s) and batching (batch size 2).
Training.
Unless otherwise stated, the DAC codec Kumar et al. (2023) and CLAP text encoder Wu et al. (2023) remain frozen; we train the FiLM-conditioned transformer masker and the lightweight query network end-to-end with an Adam optimizer Kingma and Ba (2017) and a plateau-based learning-rate schedule Mukherjee et al. (2019). We produce two variants of CodecSep, trained separately on dnr-v2 and on AudioCaps, to study distributional effects; we denote them with the suffixes “+dnr-v2” and “+AudioCaps”. For fair comparison, TDANet and CodecFormer are re-trained from scratch on dnr-v2 using our setup, AudioSep is evaluated both as a public checkpoint and when re-trained under matched protocols, and SDCodec is used as released by the authors. To reflect realistic deployments where signals traverse compression pipelines, inputs to non-codec baselines are passed through the codec during both training and inference. Full hyper-parameters, iteration schedules, batch configurations, and hardware details are deferred to the Appendix C.
@@
Model Metric ($\uparrow$) Music Speech Sfx
@@
CodecSep + dnr-v2 SI-SDR $0.16^{\pm 3.55}$ $8.26^{\pm 2.60}$ $0.96^{\pm 4.20}$
(codes in : codes out, zero-shot) ViSQOL $2.52^{\pm 0.52}$ $2.97^{\pm 0.44}$ $2.33^{\pm 0.67}$
@@ 4.1 Results and Discussions
Tables 15 present universal source separation results, prompt–granularity analyses, architectural ablations, cross-dataset generalization, paraphrase robustness, and full inference complexity under matched training/evaluation protocols. Table 1 reports dnr-v2 test results for speech, music, and SFX: text-guided models use generic prompts for speech/music and ground-truth compositional captions for SFX; we also include a masker ablation to isolate the role of the Transformer masker. Table 2 studies SFX prompt granularity across three regimes—(i) fixed-stem, non-text baselines (TDANet, CodecFormer, SDCodec), (ii) generic 3-stem prompts ({music, speech, sfx}), and (iii) a universal setup with fine-grained, compositional SFX prompts—thereby aligning input conditions for fair comparison with fixed-head systems. Table 3 isolates architecture: decoder-style generation (CodecFormer) vs. an unguided 3-stem masker variant and its text-guided counterpart, all operating in the codec latent space. To assess out-of-domain generalization, Table 4 benchmarks on five additional open-domain corpora (ESC-50, Clotho-v2, AudioSet, VGGSound, AudioCaps) with mixtures of three randomly sampled sources and prompts drawn from captions (not tied to fixed labels).. Table 5 probes prompt paraphrasing robustness by replacing the generic speech/music prompts with unseen synonymic variants at inference (zero-shot paraphrase test). Finally, Table 6 compares end-to-end and architecture-only GMACs for spectrogram-domain separation versus codec-latent masking, including the practical code-stream case. All models are evaluated against the original (uncompressed) ground truth; our methods are highlighted in bold; and we report mean and standard deviation ($1\sigma$).
Source Separation Performance on dnr-v2 (cf. Table 1). CodecSep+dnr-v2 outperforms both pretrained AudioSep (zero-shot) and retrained AudioSep+dnr-v2 across all categories, with sizable SI-SDR gains in speech (9.97 vs. 4.92/7.68 dB), music (1.15 vs. 2.46-2.46/5.55-5.55 dB), and SFX (0.89 vs. 0.34-0.34/4.66-4.66 dB). In ViSQOL, CodecSep matches or exceeds AudioSep in speech and music while slightly trailing on SFX, likely reflecting differences in SFX prompt distributions (AudioSep’s diverse training vs. CodecSep’s compositional SFX prompts from dnr-v2). Importantly, our bitstream-native variant—CodecSep+dnr-v2 (codes in: codes out, zero-shot)—evaluates the CodecSep+dnr-v2 masker directly on reconstructed embeddings $E$ from code streams (§3.2, para. 6677) without any fine-tuning: it incurs a modest drop relative to the continuous-latent path (about $1{\text{–}}2$ dB SI-SDR across sources; small ViSQOL deltas for music/speech and parity on SFX), yet still surpasses AudioSep+dnr-v2 on SI-SDR for all sources (e.g., music: 0.16-0.16 vs. 5.55-5.55 dB; speech: $8.26$ vs. $7.68$ dB; SFX: 0.96-0.96 vs. 4.66-4.66 dB). Compared to pretrained AudioSep (zero-shot), the codes-in:codes-out variant improves SI-SDR on speech/music but lags on SFX SI-SDR and ViSQOL. These results indicate that a deployment-friendly, no-finetuning bitstream path is already competitive; as discussed in §3.2, para. $6$ , light fine-tuning on $E$ or an embedding-consistency loss can close the residual gap. As part of our ablation, the lightweight CodecSep+dnr-v2 (ablate Masker) removes the transformer masker and applies FiLM directly to the encoder; it attains SI-SDR comparable to AudioSep+dnr-v2 with better perceptual speech quality, but overall separation quality drops due to FiLM perturbing the mixture latents held by the encoder.
@@
Effect of SFX Prompt Granularity during training. We evaluate three regimes: (i) fixed-stem baselines without text guidance (TDANet, CodecFormer, SDCodec) (cf. Table 2), (ii) generic 3-stem prompts (“music/speech/sfx”) (cf. Table 2), and (iii) universal prompting that retains generic cues for speech/music but uses fine-grained, compositional SFX descriptions (cf. Table 1). For (ii) and (iii), we train and evaluate separate versions of both CodecSep and AudioSep on each prompt setting. Across settings, CodecSep is competitive with or stronger than fixed-stem baselines on SFX while maintaining comparable speech/music quality (cf. Table 2). Under matched generic-prompt training/evaluation, CodecSep remains robust and surpasses spectrogram-domain AudioSep, indicating that effective separation does not hinge on carefully crafted prompts (cf. Table  1– 2). Moreover, replacing the “sfx” label with detailed SFX prompts consistently sharpens SFX extraction and, importantly, improves speech and music stems as well—suggesting that richer SFX supervision enhances overall scene disentanglement (cf. Table 1). While these controlled studies cover multiple prompt granularities, we expect additional gains from training on larger, more diverse corpora with a spectrum of prompt specificities, which we leave for future work.
@@
Model Metric ($\uparrow$) Music Speech Sfx
@@
Why use a Transformer masker instead of a decoder?(cf. Table 3). We compare (i) CodecFormer (decoder-style source generation), (ii) CodecSep (unguided, 3-stem) which uses the CodecFormer Transformer as a masker over codec latents, and (iii) CodecSep (text-guided). The results on dnr-v2-test exhibit two clear trends. First, replacing decoder-style generation with masking consistently strengthens separation across music, speech, and SFX. This aligns with our design rationale: in the DAC latent domain, the masker modulates existing, semantically structured content instead of synthesizing new signals, which (a) reduces artifacts and cross-talk leakage, (b) preserves long-range periodicity/timbre and transient organization already encoded by the codec, and (c) stabilizes optimization compared to end-to-end generation. Second, adding text guidance yields a further uniform improvement. The masker formulation concentrates Transformer capacity on selection (“where/how much” to pass) rather than generation (“what” to produce).
@@
Model Metric ($\uparrow$) Music Speech Sfx